# Yuyang Tao, Shufei Ge

### v1

@@ 4 Application
In this section, we apply our proposed method to a real dataset. The Covid-19 pandemic poses an enormous threat to global health and economic. To combat the epidemic, one of the fundamental tasks is identifying and monitoring the mutation of the SARS-COV-2 coronavirus. Studying virus evolution helps understand the basic biological characteristics, develop vaccines and drugs, and forecast trends. A traditional way to study viral mutations is constructing phylogenetic trees based on genetic sequences. However, tree structure representations can only show vertical processes. Whereas, some processes are horizontal. For virus, the homologous recombination and reassortment are typical examples lead to non-tree-like representation chanTopologyViralEvolution2013 . We apply the D-Mapper algorithm on a SARS-COV-2 coronavirus RNA sequences dataset, serving as an alternative tool to investigate both the vertical and horizontal processes of viral mutations. The dataset contains $357$ SARS-COV-2 coronavirus RNA sequences in China with different lineages from GenBase (https://ngdc.cncb.ac.cn/genbase/) in National Genomics Data Center cncb-genbase2023 .
@@
We first compute the distance matrix of input sequences, several methods are available to compute distance between RNA or DNA sequences, such as alignment algorithm based or likelihood based method, however these methods are computational expensive for large whole DNA or RNA sequences dataset. To simplify computation, we use $k$-mers frequency vector as a representation of every sequence and compute pairwised distance base on these vectors. We choose $k=3$, then every sequence is a $64$ dimension vector. Since these RNA sequences have high similarities, we perform a min-max scaling on them. Finally, we calculate the mean of each row as the filtered data.
@@
The number of intervals are set to $n=15,\alpha=0.01$ via grid tunning. The DBSCAN with a radius of $0.6$ and a minimum of samples $3$ is chosen in the clustering algorithm. The results are shown in Figure 8. The output structure represents the evolution process of these lineages. The loop structure represents the horizontal evolution process. The lineages of the two isolated nodes may indicate significant differences from others, these lineages may warrant further investigation.
Some important lineages in these nodes are shown in Table 4. The bottom green loop nodes contain many XBB variants, which from the recombination of two subvariants and cause large-scale infections worldwide VirologicalCharacteristicsSARSCoV22023 . The recombination of virus is a typically horizontal evolution process and the D-Mapper algorithm can provide some insights of these rapid and complex evolution process.
@@
Figure 8: The result of D-Mapper on SARS-COV-2 dataset.
@@ 5 Conclusion
In this paper, we propose a distribution-guided Mapper algorithm (D-Mapper) to relax the fixed intervals (or fixed overlap ratios) restriction in the classic Mapper algorithm. Our proposed algorithm combines a mixture model and the Mapper algorithm to obtain irregular intervals based on the density of projected data. With these irregular intervals, the D-Mapper algorithm can gain deeper topological insights and enhance clustering outcome. To validate the effectiveness of our proposed algorithm, we introduce the $SC_{adj}$ score to combine $SC$ scores and extended persistence diagram as a metric to reflect the performance of overlap clustering and persistence homology. We also conduct numerical experiments with different complexity to evaluate the D-Mapper algorithm by the $SC_{adj}$. The D-Mapper outperforms the classical Mapper algorithm under various parameter settings in all experiments. We also apply the D-Mapper algorithm to the SARS-COV-2 coronavirus RNA sequences, and the result shows that the D-Mapper algorithm can reflect both the vertical and horizontal evolution processes.
Our proposed algorithm is a probabilistic model-based approach that uses data intrinsic characteristics and probability models to generate distribution-guided covers and improved topological features. It is a viable alternative to non-probabilistic approaches. With the D-Mapper, once we get the distribution of the projected data, and we further can get the intervals from each distribution by quantile. Theoretically, the distribution of original data can be computed by transformation of distributions if the filter function is well proposed, this may be helpful for further theoretical analysis.
The selection of optimal number of components remains a difficult problem to address. One possible solution is to use non-parametric mixture model, such as the Dirichlet process, by introducing the Dirichlet process prior to adaptively choose the number of components. Another approach is to apply the information criteria like AIC or BIC to find a proper number of components for the mixture model.
@@
[…]
Green isolated node FL.15, EP.1, FL.2.2.1, FY.3.3, XBF.7, BQ.1.2
Bottom green loop nodes BA.5.1.3, XBB.1.9.2, XBB.1.42, XBB.1.16.7, XBB.1.17.1
Middle blue loop nodes BA.2.86, FY.1.1, XBB.1.12, XBB.1.5.32, FL.2.3.1, FL.10.1
Upper purple loop nodes EG.4.2, EG.5.2, FY.2, FL.1, XBB.2.3.8
[…]
6 Competing interests
No competing interest is declared.
7 Acknowledgments
This project was supported by the Shanghai Science and Technology Program (No. 21010502500), the startup fund of ShanghaiTech University.

### v2

@@ 3.4 Covid-19 dataset
In this section, we apply our proposed method to a real dataset. The Covid-19 pandemic poses an enormous threat to global health and economics. To combat the epidemic, one of the fundamental tasks is identifying and monitoring the mutation of the SARS-COV-2 coronavirus. Studying virus evolution helps understand the basic biological characteristics, develop vaccines and drugs, and forecast trends. A traditional way to study viral mutations is to construct phylogenetic trees based on genetic sequences. However, tree structure representations focus on vertical processes. Whereas, some processes are horizontal. For viruses, homologous recombination and reassortment are typical ways that can lead to non-tree-like representation [24]. We employed the D-Mapper algorithm to analyze a dataset of SARS-CoV-2 coronavirus RNA sequences, utilizing it as a rapid exploratory data analysis tool. This approach allowed us to examine viral mutations’ vertical and horizontal evolutionary processes across the entire genomic scale. The dataset contains $357$ SARS-COV-2 coronavirus whole RNA sequences with different lineages from GenBase (https://ngdc.cncb.ac.cn/genbase/) in National Genomics Data Center [25].
@@
We first compute the distance matrix of input sequences. There are several methods available to measure the distance between RNA or DNA sequences, such as alignment algorithm-based [26] or likelihood-based method[27], however these methods are computational expensive for large whole DNA or RNA sequences dataset [28]. To simplify computation, we use $K$-mers [29] frequency vector as a representation of every sequence and compute pairwise distance based on these vectors. We set $k=3$, then every sequence is a $64$ dimension vector. Due to the high similarity of these RNA sequences, a min-max scaling is performed on the distance matrix. The mean value function is selected as the filter function.
@@
The number of intervals are set to $n=15,\alpha=0.006$ via grid tuning in the D-Mapper algorithm. The DBSCAN with a radius of $0.6$ and a minimum of samples $3$ is chosen in the clustering algorithm. The results are shown in Figure 8, and the full list is provided in the supplementary. The output structure represents the evolution process of these lineages. The loop structure represents the horizontal evolution process. The lineages of the two isolated nodes may indicate significant differences from others, these lineages may warrant further investigation. The result of classic Mapper algorithm is also provided in the appendix.
@@
Figure 8: The result of D-Mapper on SARS-COV-2 dataset. There are $19$ nodes in total, indexed from 00 to $18$.
Some important lineages in these nodes are shown in Table 4. The bottom green loop nodes (node 9, 10, 12) contain many XBB variants, which form the recombination of two subvariants and cause large-scale infections worldwide [30]. Nodes 4 and 5 contain the JN.1 variant, which is currently widespread worldwide [31]. When compared to its predecessor, BA.2.86, the JN.1 variant only has one additional mutation within its spike protein. Nonetheless, BA.2.86 descent while JN.1 rapidly become predominant [32]. The D-Mapper graph can give some insights into this phenomenon, where BA.2.86 and JN.1 are both in a loop structure (nodes 5, 6, 8), which indicates the existence of complex horizontal evolution between these two variants. Moreover, JN.1 variant is located in both linear (node 4) and cyclical structures (node 5), suggesting it may play a pivotal role in the evolutionary process. Consequently, other variants presented at these nodes (nodes 4, 5) warrant further scrutiny. The upper purple loop (nodes 0, 1, 2, 3) encompasses variant BA.2, which is the predecessor of BA.2.86. The spike protein of BA.2.86 has more than 30 mutations than BA.2 [33], which are reflected as significant distances on the D-Mapper graph. In summary, the D-Mapper algorithm offers crucial insights into the rapid and complex evolution process of viruses.
Nodes Lineages
7, 11, 14, 16, 18 DZ.1, EP.1, FL.2.2.1, XBF.7, BN.1.1, EJ.2, DY.2
9, 10, 12 BA.5.1.3, XBB.1.9.2, XBB.1.42, XBB.1.16.7, XBB.1.17.1
5, 6, 8 BA.2.86, FY.1.1, XBB.1.12, XBB.1.5.32, FL.2.3.1, JN.1
0, 1, 2, 3 BA.2, EG.4.2, EG.5.2, FY.2, FL.1, XBB.2.3.8
[…]
4 Discussion
Our proposed algorithm is a probabilistic model-based approach that uses data intrinsic characteristics and probability models to generate distribution-guided covers and improved topological features. It is a viable alternative to non-probabilistic approaches. With the D-Mapper, once we get the distribution of the projected data, we further can get the intervals from each distribution. Theoretically, the distribution of original data can be computed by transformation of distributions if the filter function is well proposed, this may be helpful for further theoretical analysis.
The selection of the optimal number of components remains a difficult problem to address. One possible solution is to use a non-parametric mixture model, such as the Dirichlet process, by introducing the Dirichlet process prior to choose the number of components adaptively. Another approach is to apply the information criteria like AIC or BIC to find a proper number of components for the mixture model.
Although our proposed metric $SC_{adj}$ is more reasonable than $SC_{norm}$, it still has some limitations. The $TSR$ primarily assesses the stability of the Mapper graph but does not account for other aspects of the Mapper graph, for example, the richness of the topological structures. Moreover, domain knowledge is essential for interpreting a Mapper graph. It is advisable to integrate domain expertise when assessing different Mapper graphs in real applications.
@@ 5 Conclusion
In this paper, we propose a distribution-guided Mapper algorithm (D-Mapper) to relax the fixed intervals and overlapping ratios restriction in the classic Mapper algorithm. Our proposed algorithm combines a mixture model and the Mapper algorithm to obtain irregular intervals based on the density of the projected data. With these irregular intervals, the D-Mapper algorithm can gain deeper topological insights and enhance clustering outcomes. To validate the effectiveness of our proposed algorithm, we introduce the $SC_{adj}$ score to combine $SC$ scores and extended persistence diagram as a metric to reflect the performance of overlap clustering and persistent homology.
We also conduct simulation studies with different complexity to evaluate the D-Mapper algorithm by the $SC_{adj}$. In all simulations, the D-Mapper outperforms the classic Mapper algorithm concerning on the metric $SC_{adj}$. The $SC_{norm}$ of the D-Mapper are all higher than the classic Mapper algorithm, and all $TSR$s are $1$. This indicates that the D-Mapper can achieve better clustering than the classic Mapper algorithm while outputting high-quality Reeb graph approximations. Note that we tune parameters by grid searching the highest $SC_{adj}$, thus the $TSR$ is relatively high. In many cases, the $TSR$ could be low, as we show in Table 1 and Table 2. Our experimental results also indicate that our proposed metric $SC_{adj}$ is more stable than the metric $SC$. We also apply the D-Mapper algorithm to the SARS-COV-2 coronavirus RNA sequences, and the result shows that the D-Mapper algorithm can reflect both the vertical and horizontal evolution processes.
@@
Acknowledgements This project was supported by the Shanghai Science and Technology Program (No. 21010502500), the startup fund of ShanghaiTech University.
Appendix A Output of the classic Mapper algorithm on Covid-19 dataset
The result of the classic Mapper algorithm is shown in Figure 9, and the full list is provided in the supplementary.
Figure 9: The result of classic Mapper on SARS-COV-2 dataset ($n=15,p=0.49$). There are $20$ nodes in total, indexed from 00 to $19$.
Appendix B Bottleneck bootstrap algorithm
The confidence sets for extended persistence diagrams allow us to separate topological signals from noises. One effective method to compute confidence sets is the bottleneck bootstrap algorithm [20]. The bottleneck distance is often used as a metric to compute the distance between persistence diagrams. For an extended persistence diagram, points that have a short lifetime are usually considered as a noise and a long lifetime could be a signal.
Let the dataset with filtered values be a set of pairs $(X_{1},Y_{1}),...,(X_{n},Y_{n})$. The Mapper graph is represented by $M$ and the bootstrap samples are denoted as $(X^{*}_{1},Y^{*}_{1}),...,(X^{*}_{n},Y^{*}_{n})$. The Mapper graph computed from these bootstrap samples is denoted by $M^{*}$. Given a confidence level $\epsilon$, this algorithm outputs a distance $d_{\epsilon}$ which can be utilized to differentiate between noise and signal in the extended persistence diagram.
Algorithm 4 Bottleneck bootstrap
1:Input: Dataset and filtered value $(X_{1},Y_{1}),...,(X_{n},Y_{n})$, all parameters for the Mapper or D-Mapper algorithm, confidence level $\epsilon$.
2:Computing Mapper graph $M$ by the Mapper or D-Mapper algorithm by using original samples.
3:for $i=1$ to N do
4:     Bootstrap sampling: $(X^{*}_{1},Y^{*}_{1}),...,(X^{*}_{n},Y^{*}_{n})$.
5:     Computing bootstrap Mapper graph $M^{*}$ by the Mapper or D-Mapper algorithm.
6:     $d_{i}=d(M,M^{*})$
7:end for
8:Find $\alpha$ quantile from $\{d_{1},...,d_{N}\}$, denote as $d_{\epsilon}$.
9:Return $d_{\epsilon}$.